OpenStudy (anonymous):

why would the perimeter of two similar octagons with the areas of 112in^2 and 63in^2 be 4/3??

Directrix (directrix):

If two polygons are similar, then the square of the scale factor of the two polygons is the ratio of the areas of the polygons. The scale factor of the polygons is the same as the ratio of the perimeters. Let's check: (112/63) = (4/3)² ? (7*16)/(7*9) = 16/9 = (4/3)²